OpenStudy (readergirl12):

Could some one check my answer? What is the value of N? 12+4n/3 =8 n=3?

OpenStudy (anonymous):

Is this \[\dfrac{12+4n}{3}=8\]or\[12+\dfrac{4n}{3}=8\] Because what you've written suggests that it's the second equation, but your answer is correct if it's the first one.

OpenStudy (readergirl12):

It's the first one.

OpenStudy (dayakar):

if it is the first one your answer is right